OpenStudy (anonymous):

Trig identities: finding the exact value of expression. if the expression is not valid, state why it is invalid. without using calculator sin[sin^1(2)]

OpenStudy (swag):

im sure what the question is asking but the answer to the problem itself is 0.7891 i believe

OpenStudy (swag):

not sure*

OpenStudy (anonymous):

thats what i think

OpenStudy (swag):

sorry man

OpenStudy (anonymous):

:(

OpenStudy (swag):

Sorry i tried

OpenStudy (anonymous):

thanks anyways :)

OpenStudy (callisto):

you mean \[\sin (\sin2) ?\]

OpenStudy (anonymous):

no it is a trick question

OpenStudy (anonymous):

range of sine is \[[-1,1]\] so domain of arcsine is \[[-1,1]\] and therefore you cannot take the arcsine of 2 since it is not in the domain

OpenStudy (anonymous):

in other words there is no such number x with \[\sin(x)=2\] and therefore \[\sin^{-1}(2)\] does not exist

OpenStudy (anonymous):

no such number

OpenStudy (anonymous):

since \[\sin^{-1}(2)\] does not exist, neither does \[\sin(\sin^{-1}(2))\] for that matter neither does \[(\sin^{-1}(2))^2\] or any other expression containing \[\sin^{-1}(2)\]

OpenStudy (anonymous):

they are trying to trick you in to saying \[\sin(\sin^{-1}(2))=2\] reasoning that \[f(f^{-1}(x))=x\] but don't fall for the trap

OpenStudy (anonymous):

:D thank you..i have more question by the way..which ask the same valid or invalid query

OpenStudy (anonymous):

cos^-1[cos (7pi/3)]

OpenStudy (turingtest):

since \(\cos(\frac{7\pi}6)=-\frac12\) you should be able to answer the question with what satellite has said

OpenStudy (turingtest):

actually \(\cos(\frac{7\pi}6)=-\frac{\sqrt3}2\), but same answer...

OpenStudy (anonymous):

7pi/3 ..not 7pi/6 @TuringTest

OpenStudy (turingtest):

whatever, you are missing the point: \(\sin\theta,\cos\theta\le1\) for all possible \(\theta\) therefor \(\cos(\frac{7\pi}3)<1\) (it happens that \(\cos(\frac{7\pi}3)=\frac12\), but all that matters is that it is less than one) the idea here being that they want you to recognize that \(\sin^{-1}x\) and \(\cos^{-1}x\) are undefined for \(x>1\) so as long as that does not happen, the expression has meaning; since \(\cos\theta\le1\) then \(\cos^{-1}(\cos\theta)\) is defined, and so in this case we do in fact have \(\cos^{-1}(\cos\theta)=\theta\)

OpenStudy (radar):

I trust those answers provided by TuringTest and satellite73. Stay within the domain of those functions.

OpenStudy (radar):

@gbluedinosaur, unable to respond via the message feature.